OpenStudy (anonymous):

Find max value of f(x)=x raised to 1/x and f(x) = (1/x) raised to x

OpenStudy (anonymous):

for fx= x^(1/x) max value is e^(1/e) for fx= (1/x)^x max value is again e^(1/e)